OpenStudy (anonymous):

A test consists of multiple choice questions worth 3 points and essay questions worth 5 points. The test has 17 questions that are worth a total of 65 points. Use a model to determine how many multiple choice questions are on the test

OpenStudy (candyme):

x+y=17 3x+5y=65 -3x-3y=-51 3x+5y=65 2y=14 y=7 x+7=17 x=10 10 multiple choice and 7 essay :)